Machine Design

Given data for bracket is P=20kN, e=100mm, S(yt)=400N/mm², FS=3. Find the size of the bolt using maximum principal stress theory.
Find the direct shear stress in bolts if diameter=18mm if data for bracket is P=20kN
a) None of the listed
b) 19.2mm
c) 15.3mm
d) 21.2mm
Answer #1

Answer: b
Explanation: σ(max)=σ/2 + √[(σ/2)²+ τ²]=10154.35/A = 35N/mm².

answered by: anonymous